A line is perpendicular to the line y=2x +3 and has the same x-intercept as x+3y+10=0
Find the equation of this line. Express your answer in the form y=mx+b

Answer by pragya.ban@gmail.com(32)

A line is perpendicular to the line y=2x +3 and has the same x-intercept as x+3y+10=0
Find the equation of this line. Express your answer in the form y=mx+b

To make things similar, let's change x+3y+10=0 to y=mx+b form:

x+3y=-10
3y=-x-10
y=-x/3-10/3

Now we want the line to be perpendicular to y=2x+3. Perpendicular lines have negative reciprocol slopes of each other. So for the line we want, the slope has to be -1/2. We also know that the y intercept has to be -10/3 since that is what it told us. So the equation should be:

y=-1/2x-10/3

Hope this helps!